OpenStudy (anonymous):

Factor x²-7x+12 Can you show me how to do this?

Directrix (directrix):

Factor x²-7x+12 --------------- What are two numbers that multiply to 12 and add to -7 ?

OpenStudy (anonymous):

The options are a. (x-3)(x+4) b.(x-3)(x-4) c.(x-6)(x-2) d.(x-1)(x-12)

OpenStudy (anonymous):

*-3 and -4

Directrix (directrix):

x²-7x+12 = x^2 - 3x -4x +12

Directrix (directrix):

x^2 - 3x -4x +12 = x(x - 3) -4(x - 3)

Directrix (directrix):

x(x - 3) -4(x - 3) = (x-3) (x - 4)

OpenStudy (anonymous):

ok thanks! I'll try the rest on my own and post them to make sure i get it.

Directrix (directrix):

That's a good idea. I used "busting the b" to do the factoring. I'll try to find where I wrote up the technique and post it here for you to use as an example to help learn to factor.

OpenStudy (anonymous):

that would be great, thanks

Directrix (directrix):

"Bust the b" technique for factoring 4 x² -1x + 18 a = 4 and c = 18 and b = -1. The task on "bust the b" is to find numbers that mutiply to a*c AND add to b. In this case, numbers that multiply to 4*18 and add to -1. 14 * 18 breaks aparts into the product of prime factors : 2*2*2*3*3 Look at those five factor of 14 * 18 and try to put them in the form of 2 numbers that differ by -1. 2*2*2 and 3*3 appear to differ by 1. That would be 8 and 9. "b" has been "busted into 8 and - 9 whose sum is -1 which equals "b" in this problem. 4x - x+ 18 = 4 x² + 8 x - 9 x - 1. The "b" has been busted. Factoring by grouping comes next. 4 x² + 8 x - 9 x - 18 = 4x ( x + 2) - 9 ( x + 2 ) = --> On this step, (x + 2) is the common factor of the expression. [ (x + 2) ] ( 4x - 9 ) = ( x + 2 ) ( 4x - 9)